# Gradient-flow SDEs have unique transient population dynamics

### v2

@@ Appendix F Potentials
Each of these potentials is smoothly differentiable and hence Lipschitz. In order to satisfy the linear growth condition, $\|\nabla\Psi(x)\|\leq K(1+\|x\|)$, assumed for the existence of strong solutions to the SDE (1), we can multiply potentials by a smooth cutoff function, which is $1$ in $\|x\|\leq M$ and 0 in $\|x\|>2M$. Choosing $M=100$ for our experiments do not alter data generation compared to the raw setting.
@@ Appendix G Additional experiments
The code in this paper were adapted from two public code repositories (APPEX: https://github.com/guanton/APPEX and JKOnet: https://github.com/antonioterpin/jkonet-star). All computations are performed on a 2024 MacBook Pro with 1616GB RAM and an Apple M4 chip. The code was adapted to design the experiments, visualize the data, and interpret the results. It is available in the supplementary material.
@@ G.1 Experimental setup
Since the objective of our main experiments was to evaluate the inferential power of different Schrodinger bridge methods, we considered the methods WOT, SBIRR, and nn-APPEX (ours), and we simulated data from a range of population dynamics for each SDE, by randomizing the initial distribution over Gaussian mixtures. We perform two additional experiments, such that we use the same experimental setup, but consider two particular initial distributions: the uniform distribution over the region of interest $4,4]^{2}$ and the stationary Gibbs distribution $p_{\mathrm{eq}}$. In order to sample from the Gibbs distributions $\frac{1}{Z}\exp(-\frac{2\Psi(x)}{\sigma^{2}})$, we generate trajectories from $p_{0}=\textrm{Unif}([-4,4]^{2})$ and run $200$ steps of the SDE with step size $\Delta t=0.01$. Indeed, each of the potentials (36)-(40) satisfies the Poincaré inequality, which ensures that marginals converge exponentially quickly to the stationary Gibbs distribution [Pavliotis, 2014, Theorem 4.4].
We also perform additional experiments comparing our method, nn-APPEX, against the state-of-the-art variational method, JKOnet [Terpin et al., 2024]. We implement JKOnet with default hyperparameters, and use $100$ training epochs, noting rapid convergence of the objective. We verified in a separate experiment that running with $10000$ epochs did not meaningfully change performance. The experimental setting is the main setting over random GMM initializations, in order to test on different population dynamics. Since JKOnet also jointly estimates drift and diffusion for the gradient-flow SDE (1), we additionally report diffusivity MAE for this experiment.
@@ G.2 Results
The uniform initial distribution $p_{0}\sim\mathrm{Unif}[-4,4]^{2}$ represents an idealized benchmark [Terpin et al., 2024] while the stationary distribution $p_{0}\sim p_{\mathrm{eq}}$ represents the non-identifiable setting. Results comparing the three SB methods, WOT, SBIRR, and nn-APPEX, are summarized for the uniform initialization in Fig.˜4 and for the non-identifiable stationary initialization in Fig.˜5.
We also present results from our main experimental setting (random GMM initializations), which compare our method nn-APPEX against JKOnet in Fig.˜6. The results show that nn-APPEX accurately infers both the drift and diffusivity to high precision for all SDEs, and outperforms JKOnetexcept for the diffusivity estimate of the Bohachevsky potential. We note that the Bohachevsky potential is particularly difficult to estimate, as it has the highest magnitude, and converges to its Gibbs distribution $p_{\mathrm{eq}}$ at a much faster rate than the other SDEs, all while exhibiting highly periodic and nonlinear dependencies. For this reason, the MLE diffusion estimate by nn-APPEX likely achieves biased estimates, since the adjusted quadratic variation (17) is impacted by poor drift estimation. This is also consistent with previous results [Terpin et al., 2024], which showed that the Bohachevsky potential is particularly difficult to estimate in dimension $d=2$.
@@
Figure 5: The ability of different Schrödinger Bridge methods to infer the gradient-flow drift is evaluated across five potentials using (a) normalized absolute error (lower is better) and (b) cosine similarity (higher is better). Methods observe three marginals with the initial distribution equal to the SDE’s stationary Gibbs distribution, $p_{0}\sim p_{\mathrm{eq}}$. Aggregated box-and-whisker plots over $10$ seeds show that all methods perform similarly poorly, corroborating ˜4.1: the true gradient-flow drift is not identifiable without knowing the diffusivity.
@@
Figure 6: The ability of the variational method JKOnet and nn-APPEX (ours) to infer the gradient-flow drift is evaluated across five potentials using (a) normalized absolute error (lower is better) and (b) cosine similarity (higher is better), and their ability to infer the diffusivity is evaluated using (c) absolute error (lower is better). Both methods are given samples from three distinct marginals, such that the initial distribution is a Gaussian mixture model with randomly initialized components. The plots aggregated from $10$ seeds show that our method, $n$-APPEX achieve better estimation of the drift and diffusion in this setting, except for the Bohachevsky potential.

### v3

@@ Appendix F POTENTIALS
Each of these potentials is smoothly differentiable, and we note that they each lead to well-defined strong solutions to the SDE (1). Indeed, in order to satisfy the linear growth condition, $\|\nabla\Psi(x)\|\leq K(1+\|x\|)$, we can multiply potentials by a smooth cutoff function, which is $1$ in $\|x\|\leq M$ and 0 in $\|x\|>2M$. Choosing $M=100$ for our experiments does not alter data generation compared to the raw setting. Similarly, since the cutoff function ensures that $\Psi$ is supported in a compact subset of $\mathbb{R}^{d}$, it follows that smoothness of $\Psi$ implies that the drift $\nabla\Psi$ is Lipschitz.
@@ Appendix G ADDITIONAL EXPERIMENTS
The code in this paper was adapted from two public code repositories (APPEX: https://github.com/guanton/APPEX and JKOnet: https://github.com/antonioterpin/jkonet-star). All computations are performed on a 2024 MacBook Pro with 1616GB RAM and an Apple M4 chip. The code was adapted to design the experiments, visualize the data, and interpret the results. It is available in the supplementary material.
@@ G.1 Experimental Setup
Since the objective of our main experiments was to evaluate the inferential power of different Schrödinger Bridge methods, we considered the methods WOT, SBIRR, and nn-APPEX (ours), and we simulated data from a range of population dynamics for each SDE, by randomizing the initial distribution over Gaussian mixtures. We perform two additional experiments, such that we use the same experimental setup, but consider two particular initial distributions: the uniform distribution over the region of interest $4,4]^{2}$ and the stationary Gibbs distribution $p_{\mathrm{eq}}$. In order to sample from the Gibbs distributions $\frac{1}{Z}\exp(-\frac{2\Psi(x)}{\sigma^{2}})$, we generate trajectories from $p_{0}=\textrm{Unif}([-4,4]^{2})$ and run $200$ steps of the SDE with step size $\Delta t=0.01$. Indeed, each of the potentials (36)-(40) satisfies the Poincaré inequality, which ensures that marginals converge exponentially quickly to the stationary Gibbs distribution [Pavliotis, 2014, Theorem 4.4].
We also perform additional experiments comparing our method, nn-APPEX, against the state-of-the-art variational method, JKOnet [Terpin et al., 2024]. We implement JKOnet with default hyperparameters, and use $100$ training epochs, noting rapid convergence of the objective. We verified in a separate experiment that running with $10,000$ epochs did not meaningfully change performance. The experimental setting is the main setting over random GMM initializations, in order to test on different population dynamics. Since JKOnet also jointly estimates drift and diffusion for the gradient-flow SDE (1), we additionally report diffusivity MAE for this experiment.
@@ G.2 Results
The uniform initial distribution $p_{0}\sim\mathrm{Unif}[-4,4]^{2}$ represents an idealized benchmark [Terpin et al., 2024] while the stationary distribution $p_{0}\sim p_{\mathrm{eq}}$ represents the non-identifiable setting. Results comparing the three SB methods, WOT, SBIRR, and nn-APPEX, are summarized for the uniform initialization in Figure˜4 and for the non-identifiable stationary initialization in Figure˜5.
We also present results from our main experimental setting (random GMM initializations), which compare our method nn-APPEX against JKOnet in Figure˜6. The results show that nn-APPEX accurately infers both the drift and diffusivity to high precision for all SDEs, and outperforms JKOnetexcept for the diffusivity estimate of the Bohachevsky potential. We note that the Bohachevsky potential is particularly difficult to estimate, as it has the highest magnitude, and converges to its Gibbs distribution $p_{\mathrm{eq}}$ at a much faster rate than the other SDEs, all while exhibiting highly periodic and nonlinear dependencies. For this reason, the MLE diffusion estimate by nn-APPEX likely achieves biased estimates, since the adjusted quadratic variation (17) is impacted by poor drift estimation. This is also consistent with previous results [Terpin et al., 2024], which showed that the Bohachevsky potential is particularly difficult to estimate in dimension $d=2$.
Tables 4 and 5 provide further details about our simulated experimental results from the main text. Table 4 shows that the performance gap between SBIRR and nn-APPEX is highest in low potential areas, i.e. regions where the magnitude of the gradient field $\nabla\Psi$ is small. Further, most potentials from this benchmark have high magnitudes in a large portion of the evaluation grid, while SBIRR particularly struggles in low potential areas (see Figure˜1 and Figure˜2). Intuitively, if the potential is strong enough, then it can still be approximated even if diffusion is misspecified. To emphasize that nn-APPEX is especially useful for inference in low potential zones, Table˜4 stratifies the results based on the magnitude of the true gradient at each evaluation point. All results show that nn-APPEX outperforms SBIRR, especially in low potential zones, as one would expect from Figures 1 and 2. This is likely due to diffusion having an outsized noisy impact on the data in these regions. Moreover, Table 5 shows that the median (used in the main text) is actually a favourable evaluation metric for SBIRR compared to the mean (not shown in main text). Indeed, across all potentials, the gap in drift estimation performance between SBIRR and nn-APPEX is larger for the mean compared to the median. This occurs since SBIRR performs fairly well most of the time, but can still frequently catastrophically fail due to poor initial diffusion estimates.
@@
Table 4: Mean difference in normalized absolute drift error (and cosine similarity) between $n$-APPEX and SBIRR. Results are displayed for each potential and aggregated with respect to low potential zones (gradient magnitude is less than median) and high potential zones (gradient magnitude is higher than median). Higher numbers indicate a greater outperformance of nn-APPEX over SBIRR.
Gradient Magnitude
Potential Low High
Bohachevsky 0.293 (0.074) 0.184 (0.018)
Oakley-O’Hagan 0.612 (0.100) 0.253 (0.042)
Quadratric 1.397 (0.098) 0.654 (0.054)
Styblinski-Tang 0.128 (0.031) 0.075 (0.005)
Wavy Plateau 2.199 (0.035) 0.185 (0.012)
@@
Figure 5: The ability of different Schrödinger Bridge methods to infer the gradient-flow drift is evaluated across five potentials using (a) normalized absolute error (lower is better) and (b) cosine similarity (higher is better). Methods observe three marginals with the initial distribution equal to the SDE’s stationary Gibbs distribution, $p_{0}\sim p_{\mathrm{eq}}$. Aggregated box-and-whisker plots over $10$ seeds show that all methods perform similarly poorly, corroborating ˜3.1: the true gradient-flow drift is not identifiable without knowing the diffusivity.
@@
Figure 6: The ability of the variational method JKOnet and nn-APPEX (ours) to infer the gradient-flow drift is evaluated across five potentials using (a) normalized absolute error (lower is better) and (b) cosine similarity (higher is better), and their ability to infer the diffusivity is evaluated using (c) absolute error (lower is better). Both methods are given samples from three distinct marginals, such that the initial distribution is a Gaussian mixture model with randomly initialized components. The plots aggregated from $10$ seeds show that our method, nn-APPEX, achieves better estimation of the drift and diffusion in this setting, except for the Bohachevsky potential.
Table 5: Performance comparison of SBIRR and nn-APPEX methods across various benchmark potential functions, reporting mean and median error metrics.
Potential Method Mean Median
Bohachevsky SBIRR 0.9135 0.5804
$n$-APPEX 0.6749 0.4561
Oakley-O’Hagan SBIRR 0.5783 0.1786
$n$-APPEX 0.1463 0.0839
Quadratic SBIRR 1.1515 0.1911
$n$-APPEX 0.1261 0.0739
Styblinski-Tang SBIRR 0.5141 0.2245
$n$-APPEX 0.4126 0.1509
Wavy Plateau SBIRR 2.1761 0.2107
$n$-APPEX 0.9842 0.1960
G.3 Comparing APPEX and nn-APPEX
As a final experiment, we compare APPEX and nn-APPEX on a linear SDE problem in order to ensure nn-APPEX is fully expressive and can compete with APPEX in the linear domain for which APPEX was designed. Results are displayed in Table 6. As expected, APPEX does quite a bit better than nn-APPEX in terms of drift estimation since we have a closed form MLE estimate. APPEX’s normalized MAE score is significantly better, while their cosine similarities are comparable, pointing to the fact that nn-APPEX may do poorly in low drift regions. Interestingly, nn-APPEX does slightly better in terms of diffusion estimation ($0.055$ vs $0.074$).
Table 6: Inference results for $n$-APPEX and APPEX given the quadratic potential setting (see Appendix F). Both methods are given three marginals with Gaussian mixture model initialization.
Method Normalized MAE Cosine Similarity Diffusivity error
$n$-APPEX 0.1550 0.9925 0.055
APPEX (linear) 0.0135 0.9996 0.074